OpenStudy (luigi0210):

Differentiation:

OpenStudy (luigi0210):

Let \(y=sec^{-1}\) x. Find \(y'=\frac{d}{dx}~sec^{-1}~x\).

OpenStudy (anonymous):

first y=cosx right

OpenStudy (anonymous):

ok secy =x right?

OpenStudy (anonymous):

then secy tany *dy/dx=1

OpenStudy (akashdeepdeb):

y = sec^-1 x sec y = x 1/cos y = x cos y = 1/x y = cos^-1 (1/x)

OpenStudy (akashdeepdeb):

How is y = cos x? O.O

OpenStudy (anonymous):

rule costheta=1/sectheta

OpenStudy (akashdeepdeb):

Yes, I am aware of that. But this is an inverse function whose derivative has to be found. It is not saying (sec^-1 x = 1/cos x)

OpenStudy (anonymous):

ok sorry :)

OpenStudy (akashdeepdeb):

I am not sure how to do this question, truth be told. :) @Luigi0210, you should ask @satellite73 or @ganeshie.

OpenStudy (luigi0210):

Thanks for trying >.< We're suppose to use a right triangle to go along with this .-.

OpenStudy (luigi0210):

@whpalmer4 @jdoe0001

OpenStudy (abb0t):

First, reverse it: \(\sf x =\color{red}{arcec(y)}\) now, differentiate with respect to \(x\)

OpenStudy (abb0t):

Can you do this now?

OpenStudy (abb0t):

inverse it* and i mean't arcsec, not arcec. Lol also NOTE: arcec = sec\(^{-1}\)(x)

OpenStudy (abb0t):

arsec*

OpenStudy (abb0t):

ugh. ARCSEC* lol

OpenStudy (shamil98):

noob ^

OpenStudy (akashdeepdeb):

Woah, calm down guys, things are heating up here.